OpenStudy (anonymous):

Whats the function for the points (1,1) (2,3) (3,9) (4,27)

OpenStudy (anonymous):

Will give medals and become fan!

OpenStudy (anonymous):

Look to see the progression and what they are what do you know about 3, 9, 27 ? Hint: factor them if possible

OpenStudy (anonymous):

They each go up by multiplying 3

OpenStudy (anonymous):

yes. but if you factor them you can see the trend better if you show the exponents.

OpenStudy (anonymous):

I don't understand?

OpenStudy (anonymous):

\[ 3^0, 3^1, 3^2, 3^3, \] what do you think the 5th element will look like?

OpenStudy (anonymous):

\[3^{4}\]

OpenStudy (anonymous):

and the 6th element? then the \[x^{th}\] element?

OpenStudy (anonymous):

\[3^{5}\]

OpenStudy (anonymous):

i dont know what \[x ^{th}\]

OpenStudy (anonymous):

S0 \[y(0) = 3^0\] \[y(1) = 3^1\] \[y(2) = 3^2\] \[y(5) = 3^5\] So what is \[y(x)= ?\]

OpenStudy (anonymous):

\[y(6)=3^{6}\]

OpenStudy (anonymous):

Good! so lets ask what is y(x)? meaning if there is a variable what does the function look like. If \[y(6) = 3^6\] the rule is whatever in the parentheses goes to the power on the three. What do you think y(x) looks like then?

OpenStudy (anonymous):

\[y(x)=3^{x}\]?

OpenStudy (anonymous):

Yes, but now the issue is that you need to connect \[y(1) = 3^0\]\[y(2) = 3^1\]\[y(3) = 3^2\]\[y(4) = 3^3\] So now what is y(x)? If the pattern is take what's in the parentheses, subtract 1 and put as the power?

OpenStudy (anonymous):

\[y(5)=3^{4}\] ?

OpenStudy (anonymous):

yes! so your progression will be \[(1,3^0), (2,3^1), (3,3^2), ..., (5,3^4), (6, 3^5), ... (x, 3^?)\] what is the question mark equal to.

OpenStudy (anonymous):

its equal to to x-1?

OpenStudy (anonymous):

Yes, so the function would be \[y=3^{x-1}\]